# ProactivePIM: Accelerating Weight-Sharing Embedding Layer with PIM for Scalable Recommendation System

### v4

@@ IV-C Subtable Mapping Strategy
TT-Rec cannot benefit from the cache when a whole CnR is performed inside a single PIM unit. The first stage of CnR in TT-Rec is GEMV on the 1st and 2nd subtable, requiring the 3rd subtable subembedding to wait until all the vector multiplication of GEMV is complete. Even when the cache is utilized, the speedup is minimal since the computation start time of the 3rd subtable remains similar. Instead, as shown in Figure 5 (f), we distribute vector multiplications of GEMV across bank groups by spreading vectors composing a 2nd subtable subembedding. In this layout, each PIM executes a single vector multiplication instead of GEMV, allowing the 3rd subtable subembedding to be computed right after it is read. The host reduces the distributed GEMV.
During preprocessing, the PIM kernel duplicates specified subtables by sending write commands with the top bit of the physical address set to 1. The memory controller detects this bit and broadcasts the command to all bank groups. Once duplication is complete, the PIM kernel allocates the remaining subtables to available spaces. During inference, the memory controller checks the top bit and adjusts the channel and bank group in the PIM instruction for duplicated subtables.
@@ V Experimental Setup
Implementation. We implement ProactivePIM on DRAMsim3 [13], a cycle-accurate memory simulator. HBM2 $\times$128 and DDR4-3200 $\times$8 were set to base memory systems for all configurations, where Table I summarizing the specifications. We compare ProactivePIM performance against RecNMP, TRiM and SPACE. Hot subembeddings are treated as hot embeddings to maximize the utilization of existing designs that exploit temporal locality. For RecNMP, we linearly scaled the speedup of RankCache, following TRiM. We generate synthetic traces of CnR/GnR from the official DLRM code using criteo datasets [3], [19]. The embedding lookup per GnR is set to 80, with embedding sizes ranging from 128B to 512B.
@@
Area Overhead and Power Consumption. To obtain the area overhead and the power consumption, bd-PIM and bg-PIM were synthesized using Synopsys Design Compiler with 45nm CMOS technology operating at 300MHz clock frequency. The corresponding area was scaled to a 20nm DRAM process, taking into account that the DRAM process is 10$\times$ less dense compared to ASIC process [4]. The power of the SRAM cache was estimated using CACTI 6.5 [15].
VI Evaluation Result
As the first step in analyzing ProactivePIM’s effectiveness, we evaluate the impact of each proposed method. We then compare ProactivePIM’s performance with prior NMP designs for the embedding layer. Afterward, we analyze the area overhead and power consumption. The cache size is set to the largest target subtable size of QR-trick and TT-Rec, 100KB.
@@ VI-A Performance Impact of Design Optimizations
We evaluate the impact of our design choices on the speedup. Figure 7 shows speedup and energy efficiency improvements as design optimizations are incrementally applied to the ProactivePIM system. The baseline is the bd-PIM integrated with HBM2. The first two configurations, Proact_P and Proact_M, are PIM systems without SRAM cache: Proact_P adds only bg-PIM to the baseline, while Proact_M applies the subtable mapping strategy on Proact_P. Finally, Proact_T represents our complete PIM system with table-wise prefetch.
With a 512B embedding size, Proact_P and Proact_M achieve speedups of a 8% and 69% for QR-trick, and 13% and 272% for TT-Rec, respectively. The performance of Proact_P is lower than the ideal speedup target of 2×2$\times$2 × (bank groups $\times$ tccd_s/tccd_l) in HBM2, caused by the CPU-PIM communication overhead. Proact_M achieves $39\%$ performance improvement than Proact_P by eliminating the overhead, also reducing energy consumption by $32\%$. In each algorithm, Proact_T achieves 2.84×2.84$\times$2.84 × and 4.8×4.8$\times$4.8 × better results than the baseline, leveraging SRAM cache for intra-GnR locality. Proact_T outperforms in energy efficiency compared to other configurations due to decreased activation and read energy.
@@ VI-B Performance Evaluation with Previous NMP Designs
We evaluate ProactivePIM’s throughput and energy consumption performance against prior PIM systems, RecNMP, TRiM-G, and SPACE, using HBM2 as the base memory architecture with a consistent subtable mapping strategy. TensorDIMM is excluded as its configuration in HBM2 results in inefficient memory access below the granularity threshold.
ProactivePIM achieved the best speedup over all the prior works. Figure 8 (a) shows the relative performance of each previous work normalized to the TensorDIMM for each algorithm, where embedding size ranges from 128B to 512B. When 512B embedding size, Proact_T achieves 4×4$\times$4 ×, 4.7×4.7$\times$4.7 ×, 4.9×4.9$\times$4.9 × speedup compared to each configuration in TT-Rec. The result is similar to the QR trick. ProactivePIM’s superior achievement comes from efficient cache utilization by leveraging intra-GnR locality and implementing a subtable mapping strategy. The impact gets more significant for large embedding sizes due to increased data reads per bg-PIM.
Figure 8 (b) depicts the relative energy consumption of prior works and ProactivePIM, which are normalized to the result of RecNMP. Energy consumption is saved 54.1%, 53%, and 52% compared to RecNMP, TRiM-G and SPACE when mapping strategy is applied. The energy efficiency comes from the decreased memory access due to SRAM cache, along with the reduction in static energy from the inference time reduction.
Figure 8: Performance evaluation of QR-trick/TT-rec on NMP/PIM systems. (a) Speedup comparison (b) Energy consumption breakdown.
@@ VI-C Sensitivity Study of Cache Capacity
We conduct a sensitivity study on the cache capacity to prove the necessity of the table-wise prefetch method to reduce SRAM size and to show that SRAM overhead is minimal. Figure 9 shows variation in PIM area overhead when the total embedding table increases in each algorithm. As all-table prefetch needs to store all the target subtables, the area overhead gets larger as more embedding tables are employed. When the total table is 40, the area overhead of PIM on DRAM die is nearly 6% in QR-trick and 7% in TT-Rec. Although prefetch overhead exists when tTransfer is less than tPrefetch, the overhead is minimal.
@@ VII Conclusion
Acknowledgments
This work was supported in part by Samsung Electronics Company Ltd. Chae Eun Rhee is the corresponding author.

### v5

@@ IV-C Subtable Mapping Strategy
TT-Rec cannot benefit from the cache when a whole CnR is performed inside a single PIM unit. The first stage of CnR in TT-Rec is GEMV on the 1st and 2nd subtable, requiring the 3rd subtable subembedding to wait until all the vector multiplication of GEMV is complete. Even when the cache is utilized, the speedup is minimal since the computation start time of the 3rd subtable remains similar. Therefore, as shown in Figure 5 (f), we employ row-wise partitioning, distributing the vectors that form a 2nd subtable subembedding across bank groups. Each PIM performs individual vector multiplications instead of a GEMV operation, enabling the computation on 3rd subtable subembedding to be performed immediately after data is read, where only partial data is used for distributed multiplication. The host handles the final reduction.
During preprocessing, the PIM kernel duplicates specified subtables by sending write commands with the top bit of the physical address set to 1. The PIM extension detects this bit and broadcasts the command to all bank groups. Once duplication is complete, the PIM kernel allocates the remaining subtables to available spaces. During inference, the extension checks the top bit and adjusts the channel and bank group in the PIM-Inst for duplicated subtables.
@@ V Experimental Setup
Implementation. We implement ProactivePIM on DRAMsim3 [15], a cycle-accurate memory simulator. HBM2 $\times$128 and DDR4-3200 $\times$8 were set to base memory systems, where Table I summarizes the specifications. We compare ProactivePIM performance against RecNMP, TRiM and SPACE. Hot subembeddings are treated as hot embeddings to utilize existing designs that exploit temporal locality. We generate synthetic traces of CnR/GnR from the official DLRM code using criteo datasets [3], [23]. The embedding lookup per GnR is set to 80, with dimensions ranging from 128B to 512B.
@@
Area Overhead and Power Consumption. Bd-PIM and bg-PIM were synthesized using Synopsys Design Compiler with 45nm CMOS technology operating at 300MHz clock frequency to obtain the area overhead and power consumption. The area was scaled to a 20nm DRAM process, taking into account that the DRAM process is 10$\times$ less dense compared to ASIC process [4]. The SRAM cache power was estimated using CACTI 6.5 [18]. The power consumption of memory was calculated following [20].
VI Evaluation Results
As the first step in analyzing ProactivePIM’s effectiveness, we evaluate the impact of each proposed method. We then compare ProactivePIM’s performance on GnR with prior NMP designs. The cache size is set to the largest target subtable size of QR-trick and TT-Rec, 100KB.
@@ VI-A Performance Impact of Design Optimizations
We evaluate the impact of our design choices on the speedup. Figure 7 shows speedup and energy efficiency improvements as design optimizations are incrementally applied to the ProactivePIM system. The baseline is normal HBM2 with the last-level cache of the host set to 32MB. The first two configurations, Proact_P and Proact_M, are PIM systems without SRAM cache: Proact_P adds only bg-PIM to the baseline, while Proact_M applies the subtable mapping strategy on Proact_P. Finally, Proact_T represents our complete PIM system with table-wise prefetch. With a 512B embedding size, Proact_P and Proact_M achieve speedups of a 8% and 69% for QR-trick, and 13% and 272% for TT-Rec, respectively. The performance of Proact_P is lower than the ideal speedup target of 2×2$\times$2 × (bank groups $\times$ tccd_s/tccd_l) in HBM2, caused by the CPU-PIM communication overhead. Proact_M achieves $39\%$ performance improvement than Proact_P by eliminating the overhead, also reducing energy consumption by $32\%$. In each algorithm, Proact_T achieves 2.84×2.84$\times$2.84 × and 4.8×4.8$\times$4.8 × better results than the baseline, leveraging SRAM cache for intra-GnR locality. Proact_T outperforms in energy efficiency due to decreased activation and read energy. Prefetch overhead is minimal as speedup of Proact_T over Proact_M is nearly ideal.
@@ VI-B Performance Evaluation with Previous NMP Designs
We evaluate ProactivePIM’s throughput against prior PIM systems, RecNMP, TRiM-G, and SPACE, using HBM2 as the base memory architecture with a consistent subtable mapping strategy for fair comparison. TensorDIMM is excluded as its configuration in HBM2 results in inefficient memory access below the granularity threshold. Figure 8 (a) shows the relative speedup of each previous work normalized to the RecNMP for each algorithm, where embedding size ranges from 128B to 512B. When 512B embedding size, Proact_T achieves 4×4$\times$4 ×, 4.7×4.7$\times$4.7 ×, 4.9×4.9$\times$4.9 × speedup compared to each configuration in TT-Rec. The result is similar to the QR trick. ProactivePIM’s achievement comes from efficient cache utilization by leveraging intra-GnR locality and subtable mapping strategy. The impact gets more significant for large dimensions due to increased data reads per bg-PIM.
Figure 8 (b) depicts the relative energy consumption of prior works and ProactivePIM, which are normalized to the result of RecNMP. To properly measure the energy efficiency of the HBM-DIMM architecture over DIMM, two DIMMs are utilized across systems. Energy consumption is saved 54.1%, 53%, and 52% compared to RecNMP, TRiM-G and SPACE when mapping strategy is applied. The energy efficiency comes from the decreased access due to SRAM cache, along with the reduction in static energy from shorter inference time.
Figure 8: Performance evaluation of QR-trick/TT-rec on NMP/PIM systems. (a) Speedup comparison (b) Energy consumption breakdown for QR-trick.
@@ VI-C Sensitivity Study of Cache Capacity
We conduct a sensitivity study on the cache capacity to prove the necessity of the table-wise prefetch method for SRAM size reduction and to show that SRAM overhead is minimal in such method. As all-table prefetch stores all the target subtables, the area overhead gets larger as more embedding tables are employed. Figure 9 shows variation in PIM area overhead when the total embedding table increases. When the total table is 20, the area overhead of PIM on HBM2 die is nearly 17% in QR-trick and 15% in TT-Rec.